Two angles are complementary. The measure of one angle is 15* more than 0ne-half of the other. find the measure of each angle

Answer by robertb(5830)

x = one angle
90 - x = the other angle.
Then x = 0.5(90-x) +15.
x = 45 - 0.5x + 15,
1.5x = 60,
x = 40 degrees, one angle,
50 degrees, the other angle.